OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? x 2 3 4 5 6 y 32 72 128 200 288 A. 8 B. 16 C. 24 D. 40

OpenStudy (ilovecake):

2 3 4 5 6 x^2 4 9 16 25 36 y 32 72 128 200 288 y/x^2 = 8. So, 8 would be your answer which is 8.

OpenStudy (ilovecake):

Which is A I mean

563blackghost (563blackghost):

beat me to it @Ilovecake but yea its A

OpenStudy (anonymous):

can you guys help with another one @Ilovecake and @563blackghost

OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? x 2 3 4 5 6 y 24 54 96 150 216 A. 12 B. 6 C. 30 D. 18

OpenStudy (ilovecake):

I love this kind of math so yah I will help

OpenStudy (anonymous):

kk

OpenStudy (ilovecake):

Take each value of x and square it (as this is a quadratic variation). Now take each value of y and divide it by the corresponding value of x^2. Do you see a pattern?

OpenStudy (anonymous):

yes

OpenStudy (ilovecake):

okay now tell me what you think the answer is.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

18

OpenStudy (ilovecake):

hmmmm look again

OpenStudy (anonymous):

A?

OpenStudy (ilovecake):

no..... keep looking

OpenStudy (ilovecake):

24/4 = 6 54/9 = 6 96/16 = 6

OpenStudy (anonymous):

Oh...

563blackghost (563blackghost):

24 1,2,3,4,6,8,12,24 54 1,2,3,6,9,18,27,54 96 1,2,3,4,6,8,12,16,24,32,48,96 what numbers are common

OpenStudy (anonymous):

Can you help with another one?

OpenStudy (ilovecake):

SO B would be your answer.

OpenStudy (anonymous):

Variable y varies directly with x2, and y = 96 when x = 4. What is the value of y when x = 2?

OpenStudy (ilovecake):

okay

OpenStudy (ilovecake):

sure

563blackghost (563blackghost):

i agree @Ilovecake

OpenStudy (ilovecake):

:)

OpenStudy (ilovecake):

Do you have anymore? @Soniatonia

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Variable y varies directly with x2, and y = 96 when x = 4. What is the value of y when x = 2?

OpenStudy (anonymous):

wait, nvm XD thanks for your help

OpenStudy (ilovecake):

Oh okay. Good luck

563blackghost (563blackghost):

was happy to help xD

OpenStudy (ilovecake):

yeah. :)

OpenStudy (ilovecake):

Well I will show you how to o it anyway, here: if y = 96 when x = 4 and x gets split in halt and becomes 2, then split 96 in half to make 48. Easy! y = 48: